A Complete Bibliography of Publications in Almagest: International Journal for the History of Scientific Ideas

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: ~beebe/

01 April 2025 Version 1.04

Title word cross-reference

Eξoνσία [39]. και [39]. μαθηματικά [39].

İhsanoglu [239, 201, 161].

1-14 [190]. 1-2 [194]. 141 [72]. 15th [61]. 1668 [159]. 16th [90]. 17th [62, 133, 41, 57, 22, 9, 74]. 18e [110]. 18th [90, 237, 219, 112, 133, 57, 70, 110, 74, 15]. 19th [170, 171, 20, 5, 36, 89, 9].

20th [20, 25, 256, 126]. 267A [91].

4 [194].

60d2 [79]. 60d2-5 [79].

7.2 [137].

978-1-032-34745-5 [239].

a-t-elle [61]. A. [247]. Abacus [111]. Abbasid [239]. Abrahamic [139]. Absence [97]. Absolute [226]. Academia [178]. Academy [58]. Accent [209]. Accord [20]. accordance [91]. according [78]. Accounts [27]. Achievement [69, 102]. acier [160]. Actual [100]. ad [162]. Adam [147]. addition [275]. Addressing [235]. Aesthetic [128]. after [233]. Against [27, 78]. Agassiz [28]. Age [229, 109]. Agriculture [4]. Aigenler [147]. air [6]. Albert [61]. Alberto [48]. Albrecht [169]. Alchemy [93]. Aleksandar [20]. Alexander [136]. Alexandre [112, 225]. Algebra [45]. allègement [237]. Allemand [85]. Almagest [72, 107, 137, 165, 177, 141]. Almageste [107]. American [157]. Amire [13]. Anachronisms [54]. analogie [259]. analogy [259]. analysis [17, 18]. analytic [179]. anciennes [107]. Ancient [71, 136, 32, 17, 107]. angle [249]. Ankara [4]. Anne [91]. anneau [196]. Annex [91]. Annexe [194]. Anomaly [134]. anthropology [248]. anti [260]. anti-realist [260]. anti-réalistes [260]. Antievolutionary [207]. Antikythera [136, 233, 67]. Antoine [62]. Antoniou [21]. Apocalypticism [141]. Apokatastasis [233]. Approach [26]. archival [15]. Aristotelian [274]. Aristotle [72]. Arnauld [254]. Arnold [126]. Art [90]. Artes [147]. Arthur [68]. Asceticism [140]. Asia [51]. Assisted [257]. Astrological [52]. Astrology [224]. astronomica [147]. Astronomical [142, 61, 196]. astronomie [267, 110, 158]. astronomique [196]. astronomiques [61]. Astronomy [224, 14, 58, 175, 267, 110, 158]. Athens [2, 89]. atmosphere [237, 237]. attempt [54]. attested [15]. aurore [218]. Australian [209]. Author [49, 59, 65, 75, 82, 94, 105, 113, 122, 129, 137, 145, 153, 164, 176, 214, 221, 240]. authoritative [181]. Authors [10, 23, 30, 42, 183, 194, 202, 264, 278, 231, 271]. aux [91]. avant [267]. Awaiting [41].

B [213]. B. [253]. Babbage [81]. Back [118, 10, 23, 30, 42, 49, 59, 65, 75, 82, 94, 105, 113, 122, 129, 137, 145, 153, 164, 176, 183, 194, 202, 214, 221, 231, 240, 271, 278, 119]. Background [116]. Balkans [112]. ballistics [47]. Banat [160]. Bar [192]. barometric [237]. barométriques [237]. based [17]. Basil [187]. Basile [187]. became [169]. become [187]. before [267]. Beginning [112]. Benjamin [150, 276]. Bergson [135]. Bernard [22]. Bessarion [141]. best [40]. Betim [48]. better [81].

Between [87, 249, 218, 133, 17, 8, 239, 190, 256, 22, 143, 141, 100, 189, 110, 178]. Beyond [182, 139]. Bhutan [51]. Bible [188]. Biblical [157]. bibliographical [180]. Bibliographie [193]. bibliography [193, 180]. Bilim [161]. biography [169, 148]. biological [248]. biologie [243, 244]. biology [243, 244]. Biomathematics [108]. Biomedicine [108]. Blay [92]. Blue [276]. board [1, 11, 24, 31, 43, 50, 60, 66, 76, 83]. Bogle [51]. Book [220, 270, 277, 162, 112, 72, 93, 163, 239, 135, 57, 39, 91, 136, 56, 58, 73, 74, 201, 161, 48, 173, 125, 92, 20, 41, 22, 40]. Books [93, 57, 174]. borders [261]. boréale [218]. Boscovich [101, 102, 97, 99, 98, 100]. Boscovichianism [96]. Bosković [143]. boson [181]. botanique [132]. botany [132]. Boundaries [44]. Boussinesq [104]. Brasileira [29]. Brazil [199]. Brazilian [29]. breakthrough [6]. Brewster [81]. Bridges [143]. bringing [81]. Britain [124]. britanniques [22]. British [22, 51]. Brudzewo [61]. Buckhardt [85]. Bukhara [68]. Bulgarian [3]. Buneman [86, 124]. Byzantine [140, 8, 38].

c [5]. C. [162]. cadran [268]. Caesarea [187]. Calcul [34]. Calculation [34]. Calculus [199, 251, 87]. Calibration [233]. Can [229]. Canales [135]. Candelabra [192]. Candélabre [192]. Cannon [13]. Capacity [12]. Cardinal [141]. Carnot [19]. Carolingian [190]. Carolingiens [190]. Cartesian [218]. cartésien [218]. cas [189, 250, 225]. Case [25, 225, 189, 250, 156]. Catastrophe [217]. Catherine [142]. Catholic [181]. ce [192]. celestial [80]. celle [250]. centuries [133, 180, 57, 9]. Century [112, 170, 90, 237, 219, 171, 20, 62, 41, 25, 61, 256, 126, 70, 110, 36, 89, 15, 22, 74]. Césarée [187]. Châteauvert [137, 127]. change [260, 263, 262]. Changed [135]. changement [260, 262]. changements [262]. changes [257, 262]. Changing [111]. Charles [68, 246]. Chemistry [93, 162]. Chemists [178]. chimica [162]. China [57, 62, 133, 151, 211, 58]. Chinese [80, 17, 58]. choice [218]. choix [218]. chrétien [187]. Christian [187]. Christine [39]. Chronological [27]. Ciel [110]. circle [32]. Circulation [133]. Civilization [70]. Clarissa [169]. clepsydra [33]. Close [167]. coefficient [228]. Coimbra [147]. Colégio [147]. Collection [56]. Combination [101]. comet [159]. comme [250]. Comment [258]. commentaires [190]. commentaries [190]. Commentary [72]. commune [137, 127]. communication [17]. Community [212, 180]. Compendia [219]. completeness [18]. Compound [67]. Computer [257]. Concept [71, 250, 143]. Conceptual [257, 70, 111, 263, 243, 18]. conceptuelle [243]. concerning [249]. Confessor [149]. Conflict [141]. Conformément [91]. confrontations [156]. connected [35]. Conolly [68]. consideration [174]. Constantine [165]. contacts [62]. contemporaries [51]. contemporary [248]. ’Content [114, 50, 60, 66, 76, 83, 95, 106, 123, 130, 138, 146, 154, 165, 195, 203, 215, 232, 43]. Contents [31, 177, 184, 222, 241, 265, 272, 1, 11, 24]. Context [71, 64]. contexts [63]. Continuity [246]. Contribution [162]. Contributo [162]. Control [3]. Controversies [211]. Conversation [157]. Copernic [61, 267]. Copernican [267]. Copernicanism [198]. copernicienne [267]. Copernicus [61, 267, 64]. corporeity [6]. cosmetics [17]. Cosmography [140]. Cosmologists [229]. cosmology [149]. Cosmopolitan [86]. Cosmopolitanism [69, 84, 87]. Cosmos [136, 181, 149]. Costas [213]. cours [200]. Cover [120, 119]. Craft [112]. created [192]. Creation [44, 211, 190, 190]. Creationism [209, 210, 206, 212, 205]. Creationists [208]. créé [192]. critique [61, 22]. croisées [2]. Crossed [2]. current [46]. Cyrillic [112].

d [91]. d’actualité [244]. Dalsème [238]. dans [260, 192, 253, 190, 187, 225, 110, 158, 160, 262, 259]. darkness [190]. Darwin [246]. Darwinian [247]. Darwinism [163]. Date [233]. d’Athènes [2]. Davy [81]. Dawn [88]. day [192]. dealing [227]. Debate [135, 44]. debates [109]. Débats [109]. découverte [259]. Dedekind [253]. del [90, 162]. della [162]. demiurge [187, 187]. Democritus [93]. demonstration [6]. demonstrations [179]. Depiction [90]. depuis [22]. description [80]. Design [211]. destiny [40]. d’étude [250, 160]. developed [61]. Development [102, 126, 132]. développée [61]. devenue [187]. devices [33]. diagram [158]. diagramme [158]. diagrams [247]. Dial [118, 117]. Dialectical [78]. dialectics [77]. diary [159]. did [258, 61]. didactics [27]. Dieu [92]. differences [256, 259, 259]. different [18]. differential [87]. dimension [77]. disagreement [263]. discomfort [7]. Discours [192]. Discourse [192]. discoveries [81]. discovery [259]. discussions [131]. dispute [218]. dissemination [87]. Dissimulating [198]. d’Istanbul [268]. Distant [20]. Distinction [100]. Divine [9]. do [63]. Doumet [137, 127]. drag [228]. d’un [2, 250]. during [269, 180, 200, 35, 89, 109, 9, 151]. Dutch [198]. Dynamic [101]. Dynamica [255]. dynamics [126].

early [156, 275, 182, 152]. Earth [273, 110]. East [159, 17]. Eaux [2]. Échanges [22]. eclipses [158, 158]. ed [72]. editio [72]. edition [253, 253]. Editorial [43, 1, 11, 24, 31, 50, 60, 66, 76, 83]. Edler [37]. eds [20, 22]. Eduard [20]. Education [20, 238, 111, 36]. Effort [14]. Efthymios [165]. Einstein [135, 228]. Ekeland [40]. Ekmeleddin [239, 201, 161]. electricity [276]. Elementary [219]. Elements [149, 175, 74]. elle [61, 258]. Emigrant [4, 37]. Emissaries [68]. Empire [90, 103]. empiricism [244]. empirisme [244]. Encounters [167]. Encyclopedia [26]. end [41]. Engagement [155]. Engels [78, 77]. Engineering [112, 103]. engineers [21]. English [68]. enjeu [218]. Enlightenment [88, 200, 109]. enter [258]. Entre [110, 218, 190, 22, 189]. entrée [258]. envoy [51]. Epic [167]. Epigrammatic [275]. epistemological [26, 143, 19]. epistemologies [260, 260]. equation [89]. Era [182]. Erasing [261]. Ernest [16]. Errata [176]. ’Erratum [137]. especially [80]. Essay [143]. essor [132]. est [244, 258]. est-elle [258]. est-il [244]. Ethnic [235]. Étude [225, 196]. Euclid [249]. Euler [45]. Europe [133, 180]. European [171, 143, 109]. européens [109]. Evolution [71, 12, 44, 211]. evolutionary [247, 207]. Evolutionists [208]. exceptionalism [246]. exchanges [22]. Exhalation [71]. expanding [139]. expatriates [87]. expectation [168]. experiment [33]. experimentation [6]. Experimenters [170]. Explaining [55]. Exploring [9]. Exportation [209].

fabrication [160]. Faire [250]. Faith [8]. Faiths [139]. False [173]. famous [173]. Far [159]. Fate [68]. Fedos [148]. Female [140]. Feodosy [148]. Ferdinand [62, 80, 159]. Fermat [179]. Fibonacci [111]. Fiction [167, 170]. Fields [20]. figure [73]. fille [108]. finally [228]. First [199, 228, 178, 29, 201, 247, 20, 62, 38, 51, 132]. First-order [228]. Física [48]. fixed [217]. flat [228]. Fluxional [226]. foil [172]. followed [228]. fonctions [250, 110]. fondements [191, 132]. Foreward [242].

Foreword [138, 185, 96, 11]. form [190]. Formation [14, 180]. forme [190]. formes [109]. Forms [3, 109]. fort [108]. Fouad [91]. foundations [191, 132]. Foundry [13]. Four [93, 190]. Fournet [91]. Fox [22]. francais [218]. francaise [194]. franco [22]. French [243, 260, 244, 127, 191, 218, 237, 2, 85, 192, 92, 193, 189, 250, 253, 34, 255, 190, 61, 267, 187, 225, 22, 91, 188, 104, 258, 196, 268, 236, 200, 110, 132, 266, 107, 158, 109, 160, 262, 259]. Fresnel [228]. Friedrich [77]. Front [120, 1, 11, 24, 31, 43, 50, 60, 66, 76, 83, 95, 106, 114, 123, 130, 138, 146, 154, 165, 177, 184, 195, 203, 215, 222, 241, 232, 265, 272, 117]. fruits [81]. Functioning [12, 33]. functions [250, 110]. Fundamental [251]. fut [192].

gadgets [133]. Gama [159]. Game [51]. Gear [67]. Gear-Trains [67]. Geminos [274]. General [115, 116]. Generations [178]. Genèse [194, 190]. Genesis [45, 190]. Gentry [151]. Geocentric [269]. geological [48]. Geometry [238, 32, 254]. Georg [169, 267]. George [189, 41, 187, 56, 51, 141]. Georges [189, 187]. Georgios [72]. German [170, 85]. Global [209]. Gnomonic [196]. gnomonique [196]. God [92, 143]. Gök [14]. Golvers [57]. Gotha [85]. Great [41, 58, 217, 51]. Greco [224]. Greece [56, 16, 163]. Greek [41, 72, 73, 74, 150, 71, 171, 5, 38, 181, 21, 144, 35, 9, 152]. Grosseteste [191]. Guerra [162]. ’Guidelines [240, 75, 137, 176, 264, 271, 10, 23, 30, 42, 49, 59, 65, 82, 94, 105, 113, 122, 129, 145, 153, 164, 183, 194, 202, 214, 221, 231, 278]. Gulliver [171].

H. [253]. Haeckel [16]. Half [219, 20].

Haller [169]. Handbook [52]. Hanno [245]. hardback [239]. Heaven [110]. Hebræus [192]. Hegel [78]. Henri [160]. Heraclides [134]. Heritage [161]. Heron [33, 6].

Hexaamerican [189]. Hexaemera [193, 193]. hexaémérique [189]. Hexaemeron [188, 188]. Higgs [181]. High [4]. Hilbert [18]. Hindu [210]. Hipparchus [91, 188]. Hipparque [91, 188]. histoire [250]. Histoires [2]. História [48]. Historical [199, 116, 18, 131, 77, 227]. Historico [26]. historiographical [248]. historiography [261]. History [139, 27, 46, 93, 25, 56, 58, 36, 245, 162, 250, 128, 38, 54]. Holographic [3]. Holy [142]. homme [192, 92]. horizons [158]. House [239, 201]. Hu [17]. Human [9]. Humanities [20]. Humour [103]. Hunters [182].

Iannis [73]. Iconography [142]. Idea [29]. idealism [78]. Ideas [16, 173, 21]. ideological [156, 174]. ignored [276]. II [192]. Illustrations [121]. Imagination [166]. Impersonal [27]. importance [51]. Impredicativity [252]. improvement [187]. Incommensurabilité [260]. incommensurability [260, 260]. Indications [67]. Indifference [155]. Indirect [172]. inertia [227]. influence [61, 61]. Initial [233]. Inscription [120, 119]. Inscriptions [118, 115, 117]. Institute [4]. Institutions [21]. instruction [147]. Instruments [142, 133]. Integral [251]. integrals [256]. Intelligent [211]. international [180]. Interpretation [157, 237, 228, 143, 237]. Interpreter [41]. intertextual [170]. Intertheoretic [64]. Introduction [199, 186, 223, 204, 84, 166, 38, 274]. Introductory [235]. Intuition [69]. inv [91]. invaluable [275]. Ioannis [72]. Irene [268, 268]. Iron [79]. Irrational [182]. ISBN [239]. Islam [206]. Islamic [63]. issue [218, 267]. Issues [235, 227]. Istanbul [37, 268]. Itô [251]. Italian [162]. Italy [111]. Ivar [40].

J [102, 143]. Japan [207]. Jean [91, 188].

Jean-Luc [91]. Jesuit [57, 159]. Jesuits [62]. jeune [108]. Jewish [41]. Jimena [135]. Johann [85, 169]. John [188]. Joly [22]. Jones [136]. José [29]. Joseph [104]. jour [192]. Journal [103]. Jules [172, 173, 238, 175, 174].

Kairis [73]. KAM [126]. Karas [73]. Karl [85]. Kaspar [26]. Kind [167]. Kinds [262]. Kingdom [162]. Kiril [47]. Know [178]. Know-How [178]. knowledge [63, 133, 8, 9]. Kolchinsky [20]. Kolmogorov [126]. Komnenos [52]. Korea [80, 205]. Kostas [163]. Kostov [112]. Koutzakiotis [41]. Koyré [225]. Kragujevac [20]. Krasovsky [148]. Krimbas [163, 213].

l [237]. Laboratory [3]. Lagrange [266]. Lambin [108]. land [171]. Language [14, 200]. langue [200]. Laplace [7]. late [62, 140, 103]. late-17th [62]. Latin [219]. Latitude [234]. Lavoisier [162]. law [179]. Laws [101]. Legacy [69, 245]. Leibniz [255, 87, 259]. Leme [48]. Lesvios [150, 276]. letter [159]. letters [73]. leur [253]. life [81, 109, 53]. light [191]. lightening [237]. Lights [218]. linear [27]. lines [254]. Links [20]. literature [189, 275]. littérature [189]. livre [22]. local [228]. l’Occident [92]. Logical [34, 244]. Logique [34, 244]. Logos [9]. Longitude [216]. look [261, 61]. Lorentz [228]. lost [80]. Louis [28]. lu [188]. Luc [91]. Luis [159]. lumière [191]. Lumières [200, 109]. Lunar [197].

Macau [159]. machine [62]. Magicians [170]. Making [81]. man [192, 92, 149]. Maney [93]. Manolis [74]. Manuel [52]. manufacturing [160]. Mapping [56]. Maps [56]. March [159]. Margarita [56]. marine [109]. Martelli [93]. Martian [217]. Martin [37]. material [133]. materialism [78]. Mathematical [199, 147, 70, 250, 255, 180, 26, 275, 111]. mathematicians [180]. Mathematics [219, 26, 275, 40, 35, 89, 175, 152, 46, 156, 266, 39]. mathématique [250, 255]. mathématiques [250, 266]. matière [190]. Matteo [93]. Matter [10, 23, 1, 11, 30, 42, 24, 31, 59, 50, 65, 75, 60, 66, 82, 94, 76, 83, 105, 113, 95, 106, 122, 129, 114, 123, 137, 145, 130, 138, 153, 164, 146, 154, 176, 183, 165, 177, 194, 202, 195, 214, 203, 221, 231, 215, 222, 240, 241, 232, 271, 278, 265, 272, 49, 43, 184, 190]. Maximus [149]. Measure [229, 32]. Measurement [225]. Measurements [230]. Mécanique [104]. mécanisme [218]. Mechanical [101, 70, 233]. mechanics [104]. Mechanism [136, 233, 67, 218, 36]. Media [211]. mediator [149]. medieval [275]. Mediterranean [12]. Mercier [91]. Messiah [41]. mesure [225]. métamorphose [243]. metamorphoses [187, 187]. metamorphosis [243]. metaphysics [243]. métaphysique [243]. Meteorologica [72]. Météorologie [236]. Meteorology [236]. Method [28, 152]. methodological [6]. metric [228]. Michael [38]. Michel [92]. Milan [86]. milieu [9]. Mind [100]. Mirası [161]. Miscellanies [140]. Mises [37]. Mission [57, 68, 51]. Mito [245]. Model [206, 169]. Modern [182, 102, 100, 201, 38, 73]. Modernisation [151]. Modernity [207, 157]. moments [245]. monades [243]. monads [243]. Monastery [142]. monde [191, 190]. Montpellier [218]. Moon [173]. Moser [126]. Movement [151]. Mulerius [198]. Multireligious [235]. Multireligious-Ethnic [235]. Multitude [238]. Museo [90]. museums [53]. Muslim [201]. mutually [35]. Myth [239].

Naples [162]. Napoli [162]. Narratives [63]. national [81]. nationalism [87]. Natural [38, 102, 74, 5, 238]. Naturalis [97]. Nature [182, 100, 92, 78]. nautical [15]. Nekrasov [156]. Neohellenic [200]. néohélleniques [200]. network [2]. Newton [273]. Newtonian [226]. Nicolaidis [165]. Nicolaus [198]. Nikola [131, 69]. Nikolaevich [148]. Noël [57]. non [27]. non-linear [27]. Normative [8]. Northern [218]. notations [259]. note [62]. Notes [199, 235]. Notion [26]. notions [18]. nouveau [61]. Nouvelle [225]. novæ [267]. novelistic [172].

Object [152, 250]. objects [133]. objet [250]. Observation [224, 80]. Observations [116, 80, 91]. observatoire [236, 236]. observatory [236]. Occasional [99]. œuvre [187]. ongoing [81]. ont [188]. order [228]. origin [190]. originalité [92]. originality [92]. origine [190]. Orthodox [157, 181, 212]. Orthodoxy [155, 144]. Oscar [86, 124]. Osmanlı [161]. Ottoman [103, 13, 235, 35, 15, 161]. our [135]. outil [187].

P. [156, 112]. Pachymeres [72]. Pacioli [45]. Paes [48]. Page [50, 66, 76, 83, 1, 11, 24, 31, 43, 60]. Paintings [90]. Pantahekineton [150]. Pappus [249]. papyrus [91]. paradigm [274]. Paradigma [112]. parallel [80]. Parapegma [117]. Paris [218, 2, 85, 62, 236]. Parthenope [162]. participants [51]. Particle [102]. Pascal [168]. path [86]. Patiniotis [74]. Patronage [140, 124]. Peking [159, 80]. pelecinum [137, 127]. Pensées [168]. perfectionnement [187]. period [35]. periodicals [5]. Periplus [245]. Persecution [124]. perspective [17, 248]. perspectives [181]. Petersburg [25]. Petrović [20]. Peurbach [61, 267]. Pharmacy [178]. phase [29]. phenomena [274, 274]. phenomenon [80]. Phili [39]. Philippes [196]. Philippi [196]. Philo [187]. Philon [187]. Philopon [188].

Philosopher [135, 73]. Philosophia [72]. Philosophiae [97]. philosophical [181]. philosophie [244, 189, 266, 225]. Philosophy [102, 9, 74, 244, 189, 266, 225]. physical [77, 104]. Physicist [135]. physico [255]. physico-mathematical [255]. physico-mathématique [255]. Physics [5, 102, 128, 227, 258]. physique [104, 258]. Pioneer [108]. Pisides [187, 187]. Pisidia [189]. Pisidie [189]. Planetary [216, 234, 67]. Plate [118]. Plato [187, 188, 79]. Platon [187, 188]. pluie [237]. Pneumatics [33, 6]. Poetic [170]. Poetry [99, 275]. Poincaré [228, 160, 252]. Political [174]. Politics [209, 141, 208]. Polygenism [28]. Pontus [134]. Popoff [47]. Popularization [29]. Popularizing [14, 5]. portable [136]. Portuguese [180, 57]. possible [40]. Potential [100]. pour [266]. powder [17]. powers [172, 259]. pp [137]. Practical [3]. Practices [205]. Practicing [205]. Prado [90]. Prague [219]. pré [267, 267]. Pre-Copernican [267]. pré-copernicienne [267]. Precision [224, 225, 225]. Preface [115]. preferred [250]. premier [132]. Presence [97]. present [227]. Presocratic [269]. Primatology [245]. princeps [72]. Principia [273]. Principles [19, 128]. Priority [152]. private [7]. privilégié [250]. Probability [168]. Problems [235, 46]. Profession [112]. Program [269]. projects [180]. prominent [73]. proneurs [158]. Proofs [257]. proposals [46]. proposed [276]. prosneuses [158]. pseudo [268, 93]. pseudo-cadran [268]. Pseudo-Democritus [93]. pseudo-sundial [268]. Ptolémée [158]. Ptolemy [64, 158]. Publishing [93]. puissances [259]. Pure [96].

Qizheng [197, 216, 234]. Quantum [101]. quatre [190]. questions [200]. qui [192].

R [247]. R. [253, 102, 143]. rôle [259]. Race [208]. rain [237]. Rationality [3]. rattachées [193]. Raymond [91]. Re [217]. Re- [217]. read [188]. Readers [57]. realist [260]. réalistes [260]. Reality [239]. Really [229]. realm [131]. reappraisal [181]. Reassessment [28]. Reception [16]. Reduction [79]. réécritures [253]. Reflections [19, 61]. réflexions [61]. refraction [179]. regard [61]. Regno [162]. Reign [52]. related [33, 193]. Relation [102, 227]. relations [8]. Relationship [111]. Relativity [228]. Relectures [253]. relevant [244]. Religion [139, 156, 143, 208, 8]. Religious [205, 7]. remarks [249, 107]. Remarques [107]. Republic [198, 29]. Rereadings [253]. réseau [2]. Revealing [136]. Review [162, 112, 92, 72, 93, 163, 239, 135, 57, 39, 91, 136, 56, 58, 201, 161, 20, 41, 22, 40, 73, 74]. Reviews [220, 270, 277]. Revista [29]. Revolting [198]. revolution [267, 267]. rewritings [253]. Richard [37, 225]. Richardson [169]. Riemann [253]. ring [196]. rivalry [170]. Robert [191, 22]. Roberval [34]. Roemer [62]. Roger [101, 99]. role [169, 133, 87, 259]. Roman [224]. Romanian [98]. Rose [157]. rota [147]. rotation [273]. Routledge [239]. Royal [13, 236, 236]. Rule [35]. Russia [25, 110, 132]. Russian [20, 68, 178]. Russie [110, 132].

S.J. [62]. Sadi [19]. Saint [149, 268, 148, 142]. Sainte [268]. salvation [149]. Sami [14]. Samourkas [56]. Sample [48]. Samuel [169]. Sanctuaires [192]. Sanctuaries [192]. Saros [233]. Satsue [245]. savants [22]. saving [274]. Schelling [78]. schemes [247]. Scholars [235, 63, 22]. scholia [107]. Schools [209, 111]. Schott [26]. Science [139, 155, 90, 182, 27, 20, 167, 157, 235, 14, 25, 44, 238, 205, 143, 70, 208, 144, 29, 166, 261, 172, 255, 8, 38, 81, 77, 54, 36, 263, 262, 90, 9, 151]. Sciences [58, 201, 5, 38, 236, 200, 262, 38, 26]. Scientific [14, 136, 28, 3, 84, 151, 161, 260, 131, 173, 74, 262]. scientifique [260]. scientifiques [262]. Scientist [37, 143, 73]. scientists [4]. scolies [107]. sea [171]. Search [100]. Second [219]. secondaire [193]. secondary [193]. Secrets [182]. sectors [35]. sélective [193, 193]. ’Self [151]. ’Self-strengthening’ [151]. Semseddin [14]. sens [104]. sense [104]. Serbian [20]. Series [58]. sérieuse [108]. serpentine [86]. SHAQA [103]. sic [242]. siècle [237, 61, 110, 109, 22]. Similarities [256]. Simple [96]. Sinai [142]. since [22]. sixième [192]. sixth [192]. Skordoulis [165]. Sky [14]. Slavic [140]. Social [211, 132]. Socialisation [151]. sociaux [132]. societies [63]. Society [93, 111]. Sofia [112]. solaire [268]. Solar [134, 197]. Some [249, 80, 256]. son [187]. Sortes [262]. soul [81]. sources [15]. South [205]. Sovereignty [39]. Soviet [156]. Space [100]. Spanish [90, 182]. spatially [228]. speech [172]. Spiral [233]. St. [25]. Stag [157]. Standpoint [100]. Stanford [86]. State [144]. statement [1]. States [44]. Statistical [101]. statistics [258]. statistique [258]. steel [160]. Stephanides [38]. still [244]. Stochastic [251]. Stoddart [68]. storia [162]. stories [2]. story [125]. straight [254]. strategies [170]. Streams [111]. strengthening’ [151]. Structures [70]. struggle [81]. Study [199, 219, 25, 47, 225, 211, 250, 196, 160]. Studying [3]. subsequent [51]. sundial [268]. Superstition [207]. Superstitious [182]. Supplementary [121]. Survie [189]. Survival [189]. Suzanne [108]. synthetic [179].

t [61]. ’Table [232, 43, 50, 60, 66, 76, 83, 265, 1, 11, 24, 31, 95, 106, 123, 130, 138, 146, 154, 165, 177, 184, 195, 203, 215, 222, 272].

Teaching [219, 46]. Technical [178]. Technique [79, 17]. Technology [35, 15]. Telelis [72]. tell [63]. Telling [250]. temps [225]. Ten [165, 177]. ténèbres [190]. terminologie [200]. terminology [200, 15]. Terra [48]. Terre [110]. Territories [98]. Territory [29]. Tertiary [167]. Tesla [69]. testimonies [159]. text [275]. Textbook [45, 38]. textbooks [48, 8, 53]. texte [194]. textes [253]. texts [253]. textual [275]. Their [57, 63, 253]. thématiques [193]. themes [193]. Theologian [143]. théologie [189]. theology [189]. Theophilos [73]. Theorem [251]. Theoria [97]. Theoricæ [267]. Theories [197]. Theory [71, 216, 234, 207, 150, 276, 126]. Thermodynamics [19]. things [9]. Thinking [32, 78]. Thomas [62]. Thought [71, 207, 74]. Three [227, 35]. Ti [79]. Tibet [51]. Tihon [91]. Time [226, 273, 224, 230, 225, 228, 227, 135]. Times [269]. title [150, 112, 41, 72, 73, 74, 1, 11, 24, 31, 43, 50, 60, 66, 76, 83]. Today [212]. Tolias [56]. tool [187]. top [143]. Tophane [13]. Tophane-I [13]. toujours [244]. tournant [237]. Tracing [15]. Trade [35]. tradition [275]. traduction [194]. Traduire [200]. Trains [67]. Trajectory [47]. Transformations [182, 189]. transition [64]. Translating [200]. Translations [97, 171]. Transmutation [28]. travel [171]. treasury [81]. Treatise [14]. Trebizond [141]. Tree [53]. trip [160]. trisection [249]. Tuibu [197, 216, 234]. Turkish [206, 14]. turn [237, 180]. Two [245]. Tycho [217]. typology [54].

Understanding [55, 135]. unidentified [80]. Union [100]. United [44]. univers [61]. universe [61, 229]. University [201, 219, 37, 89]. unrecognized [36]. unsolved [89]. unveiling [77]. urbain [2]. urban [2]. using [26].

V [163]. vapeurs [237]. vapors [237]. Var [137, 127]. variations [237]. Varieties [155]. Vavilov [131]. Vector [199]. Verbiest [62, 159, 80]. Veríssimo [29]. Verne [172, 173, 175, 174]. versus [170]. vie [109]. Viète [249]. views [174]. Visitor [159]. volumes [58]. voyage [160]. vs [226, 81, 263]. Vuillemin [266].

Wallace [247]. War [112]. wars [47]. Wartime [207]. was [192]. waters [2, 2]. Weber [253]. West [17, 92]. Western [70, 151]. Westfall [225]. Which [78, 187]. Wind [71]. Windmills [12]. Wisdom [239]. without [230]. Women [5]. Wonder [136]. Work [143, 172, 179, 187, 175]. World [139, 112, 191, 41, 190, 77, 182, 136, 201]. worlds [40].

XIX [180]. XVe [61]. XVIIe [22]. XVIIIe [237]. XX [180].

Years [165, 177, 124]. Yiannis [21]. York [239].

Zealand [208]. Zimmermann [169].

References

Anonymous:2010:FMTa

[1] Anonymous. Front matter (“Title page”, “Editorial board”, “Table of contents”, “Editorial statement”). Almagest: International Journal for the History of Scientific Ideas, 1(1):2–5, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.05048500050400028275078408817103.

Chatzis:2010:EPE

[2] Konstantinos Chatzis and Georgia Mavrogonatou. Eaux de Paris, eaux d’Athènes, 1830–1930: Histoires croisées d’un réseau urbain. (French) [Waters of Paris, waters of Athens, 1830–1930: Crossed stories of an urban network]. Almagest: International Journal for the History of Scientific Ideas, 1(1):6–20, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMA.3.1.

Tchalakov:2010:CMC

[3] Ivan Tchalakov. Control, more control: Studying practical forms of scientific rationality in a Bulgarian holographic laboratory. Almagest: International Journal for the History of Scientific Ideas, 1(1):22–48, 2010. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.2.

Kadioglu:2010:ESA

[4] Sevtap Kadioğlu. Emigrant scientists in Ankara High Institute of Agriculture. Almagest: International Journal for the History of Scientific Ideas, 1(1):50–65, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMA.3.3.

Giannakopoulou:2010:WPP

[5] Polyxeni Giannakopoulou. Women and physics popularizing natural sciences in the 19th c. Greek periodicals. Almagest: International Journal for the History of Scientific Ideas, 1(1):66–85, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMA.3.4.

Papadopetrakis:2010:DCA

[6] Eftichis Papadopetrakis and Vaios Argyrakis. The demonstration of the corporeity of air in Heron’s Pneumatics: a methodological breakthrough in experimentation. Almagest: International Journal for the History of Scientific Ideas, 1(1):86–102, 2010. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.5.

Hahn:2010:LPR

[7] Roger Hahn. Laplace’s private religious discomfort. Almagest: International Journal for the History of Scientific Ideas, 1(1):104–110, 2010. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.6.

Katsiampoura:2010:FKN

[8] Gianna Katsiampoura. Faith or knowledge? Normative relations between religion and science in Byzantine textbooks. Almagest: International Journal for the History of Scientific Ideas, 1(1):112–122, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMA.3.7.

Vlahakis:2010:PKT

[9] George N. Vlahakis. Philosophy is the knowledge of things divine and human. Exploring the divine logos in the milieu of Greek Science during the 17th– 19th centuries. Almagest: International Journal for the History of Scientific Ideas, 1(1):124–139, 2010. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.8.

Anonymous:2010:BMGa

[10] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 1(1):140–143, 2010. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.5555/J.ALMAGEST.
                                        6.76070700058005060300780380780407.

Anonymous:2010:FMTb

[11] Anonymous. Front matter (“Title page”, “Editorial board”, “Table of contents”, “Foreword”). Almagest: International Journal for the History of Scientific Ideas, 1(2):1–7, 2010. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.5555/J.ALMAGEST.
                                        6.69050402020472050002017207050888.

Bir:2010:EFC

[12] Attila Bir and Mustafa Kaçar. Evolution, functioning and capacity of the Mediterranean windmills. Almagest: International Journal for the History of Scientific Ideas, 1(2):8–26, 2010. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMA.3.9.

Ayduz:2010:ORC

[13] Salim Ayduz. The Ottoman Royal Cannon Foundry: “Tophane-I Amire”. Almagest: International Journal for the History of Scientific Ideas, 1(2):28–39, 2010. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMA.3.10.

Kadioglu:2010:SST

[14] Sevtap Kadioglu. Şemseddin Sami’s treatise of astronomy Gök (sky): an effort towards the formation of Turkish scientific language and popularizing science. Almagest: International Journal for the History of Scientific Ideas, 1(2):40–49, 2010. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.11.

Zorlu:2010:TTT

[15] Tuncay Zorlu. Tracing technology through terminology: Ottoman nautical terminology as attested in the 18th century archival sources. Almagest: International Journal for the History of Scientific Ideas, 1(2):50–83, 2010. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.12.

Kyriakou:2010:REH

[16] Kyriakos Kyriakou and Constantine Skordoulis. The reception of Ernest Haeckel’s ideas in Greece. Almagest: International Journal for the History of Scientific Ideas, 1(2):84–103, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMA.3.13.

Huang:2010:NPA

[17] Huang Huang and Ying Qin. A new perspective on ancient technique communication of cosmetics between the East and the West, based on the analysis of Chinese cosmetics — “Hu” powder. Almagest: International Journal for the History of Scientific Ideas, 1(2):106–115, 2010. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.14.

Karela:2010:HDN

[18] Catherine Karela. Hilbert on different notions of completeness: a conceptual and historical analysis. Almagest: International Journal for the History of Scientific Ideas, 1(2):116–127, 2010. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMA.3.15.

Pisano:2010:PSC

[19] Raffaele Pisano. On principles in Sadi Carnot’s Thermodynamics (1824). Epistemological reflections. Almagest: International Journal for the History of Scientific Ideas, 1(2):128–179, 2010. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMA.3.16.

Djuric:2010:BRA

[20] Djordje Djuric. [Book review:] Aleksandar Petrović and Eduard I. Kolchinsky (eds), A Distant Accord — Russian– Serbian Links in the Fields of Science, Humanities and Education: the 19th the first half of the 20th century, Kragujevac. Almagest: International Journal for the History of Scientific Ideas, 1(2):182–184, 2010. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.3.17.

Tampakis:2010:YAG

[21] Constantine Tampakis. Yiannis Antoniou, the Greek engineers: Institutions and ideas 1900–1940. Almagest: International Journal for the History of Scientific Ideas, 1(2):184–186, 2010. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMA.3.18.

Nicolaidis:2010:CLR

[22] Efthymios Nicolaidis. [Critique de livre]: Robert Fox, Bernard Joly (eds), Échanges franco–britanniques entre savants depuis le XVIIe siècle. (French) [[Book review:] Robert Fox, Bernard Joly (eds), Franco–British exchanges between scholars since the 17th century]. Almagest: International Journal for the History of Scientific Ideas, 1(2):186–188, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMA.3.19.

Anonymous:2010:BMGb

[23] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 1(2):190–192, 2010. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.77000203080201070405020308060204.

Anonymous:2011:FMTa

[24] Anonymous. Front matter (“Title page”, “Editorial board”, “Table of contents”). Almagest: International Journal for the History of Scientific Ideas, 2(1):1–3, 2011. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.80010808070400060706020184080302.

Kolchinsky:2011:HSR

[25] Eduard I. Kolchinsky. History of science in Russia in the 20th century: St. Petersburg as a case study. Almagest: International Journal for the History of Scientific Ideas, 2(1):4–34, 2011. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.3.20.

Knobloch:2011:NMH

[26] Eberhard Knobloch. The notion of mathematics — an historico– epistemological approach using Kaspar Schott’s Encyclopedia of all mathematical sciences. Almagest: International Journal for the History of Scientific Ideas, 2(1):36–59, 2011. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.3.21.

Cornelis:2011:ACI

[27] Gustaaf C. Cornelis. Against chronological and impersonal accounts of the history of science. Towards nonlinear didactics. Almagest: International Journal for the History of Scientific Ideas, 2(1):60–75, 2011. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.3.22.

Roberts:2011:LAS

[28] Jon H. Roberts. Louis Agassiz on scientific method, polygenism, and transmutation: a reassessment. Almagest: International Journal for the History of Scientific Ideas, 2(1):76–99, 2011. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.3.23.

Vergara:2011:PSI

[29] Moema de Rezende Vergara. The popularization of science and the idea of territory in the Brazilian First Republic: the José Veríssimo phase of the Revista Brasileira (1895–1900). Almagest: International Journal for the History of Scientific Ideas, 2(1):100–120, 2011. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMA.3.24.

Anonymous:2011:BMGa

[30] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 2(1):122–128, 2011. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.07020203010105010805000701060507.

Anonymous:2011:FMTb

[31] Anonymous. Front matter (“Title page”, “Editorial board”, “Table of contents”). Almagest: International Journal for the History of Scientific Ideas, 2(2):1–3, 2011. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.87020203040004027887047678078505.

Yuste:2011:AGT

[32] Piedad Yuste. Ancient geometry: Thinking about how to measure the circle. Almagest: International Journal for the History of Scientific Ideas, 2(2):4–5, 2011. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.4.1001.

Argyrakis:2011:CEC

[33] Vaios Argyrakis. The clepsydra experiment, clepsydra’s functioning and the related devices in Heron’s Pneumatics. Almagest: International Journal for the History of Scientific Ideas, 2(2):16–27, 2011. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.4.1002.

Jullien:2011:CLR

[34] Vincent Jullien. Le calcul logique de Roberval. (French) [Roberval’s logical calculation]. Almagest: International Journal for the History of Scientific Ideas, 2(2):28–61, 2011. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.4.1003.

Terdimou:2011:MGT

[35] Maria Terdimou. Mathematics, Greek trade and technology during the period of the Ottoman Rule: three mutually connected sectors. Almagest: International Journal for the History of Scientific Ideas, 2(2):62–79, 2011. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.4.1004.

Tampakis:2011:UMH

[36] Kostas Tampakis. The unrecognized mechanism: History of science education in the 19th century. Almagest: International Journal for the History of Scientific Ideas, 2(2):80–101, 2011. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.4.1005.

Kadioglu:2011:ESI

[37] Sevtap Kadioğlu and Gaye Şahinbaş Erginöz. An emigrant scientist in Istanbul University: Richard Martin Edler von Mises (1883–1953). Almagest: International Journal for the History of Scientific Ideas, 2(2):102–123, 2011. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.4.1006.

Katsiampoura:2011:BSF

[38] Gianna Katsiampoura. The Byzantine sciences in the first modern Greek history of science textbook: Michael Stephanides, Introduction to the History of Natural Sciences. Almagest: International Journal for the History of Scientific Ideas, 2(2):124–139, 2011. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.4.1007.

Nicolaidis:2011:BRC

[39] Efthymios Nicolaidis. Book review: Christine Phili, Eξoνσία και μαθηματικά (Sovereignty and Mathematics). Almagest: International Journal for the History of Scientific Ideas, 2(2):140–145, 2011. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.4.1008.

Sheynin:2011:BRI

[40] Oscar Sheynin. [Book review:] Ivar Ekeland, The best of all possible worlds. Mathematics and destiny. Almagest: International Journal for the History of Scientific Ideas, 2 (2):146–147, 2011. CODEN ???? ISSN 1792-2593. URL https:/

/www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.4.1009.

Katsiampoura:2011:BRG

[41] Gianna Katsiampoura. [Book review:] George Koutzakiotis, [title in Greek] (Awaiting the end of the world in 17th century: the Jewish Messiah and the Great Interpreter). Almagest: International Journal for the History of Scientific Ideas, 2(2):148–149, 2011. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.4.1010.

Anonymous:2011:BMGb

[42] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 2(2):150–151, 2011. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.68030304030806030902040001050005.

Anonymous:2012:FMTa

[43] Anonymous. Front matter (‘Title page’, ‘editorial board’, ‘table of contents’). Almagest: International Journal for the History of Scientific Ideas, 3(1):1–3, 2012. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.80040474820801020002050203040704.

Larson:2012:CEB

[44] Edward J. Larson and Ronald L. Numbers. Creation, evolution, and the boundaries of science: The debate in the United States. Almagest: International Journal for the History of Scientific Ideas, 3(1):4–24, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100793.

Heeffer:2012:GAT

[45] Albrecht Heeffer. The genesis of the algebra textbook: From Pacioli to Euler. Almagest: International Journal for the History of Scientific Ideas, 3(1):26–61, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100794.

DeVittori:2012:HMT

[46] Thomas De Vittori. History in mathematics teaching: current problems and new proposals. Almagest: International Journal for the History of Scientific Ideas, 3(1):62–77, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100795.

Mazliak:2012:STK

[47] Laurent Mazliak. Study of a trajectory: Kiril Popoff, wars, and ballistics. Almagest: International Journal for the History of Scientific Ideas, 3(1):78–105, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100796.

Figueiroa:2012:SGT

[48] Silvia F. de M. Figueirôa. A sample of geological textbooks: the book História Física da Terra (1943) by Alberto Betim Paes Leme. Almagest: International Journal for the History of Scientific Ideas, 3(1):106–121, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100797.

Anonymous:2012:BMGa

[49] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 3(1):122–127, 2012. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.76030801000607040788810205010200.

Anonymous:2012:FMTb

[50] Anonymous. Front matter (‘Title page’, ‘Editorial Board’, ‘table of contents’). Almagest: International Journal for the History of Scientific Ideas, 3(2):1–3, 2012. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.82090076060202060802050006020404.

Postnikov:2012:GBF

[51] Alexey V. Postnikov. George Bogle, the first British envoy to Bhutan and Tibet (1774–1775): the importance of his mission for his contemporaries and subsequent participants of the Great Game in Asia. Almagest: International Journal for the History of Scientific Ideas, 3(2):4–45, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100822.

Caudano:2012:AHR

[52] Anne-Laurence Caudano. An astrological handbook from the reign of manuel I komnenos. Almagest: International Journal for the History of Scientific Ideas, 3(2):46–65, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100823.

Torrens:2012:TLT

[53] Erica Torrens and Ana Barahona. The ‘Tree of Life’ in textbooks and museums. Almagest: International Journal for the History of Scientific Ideas, 3(2):66–89, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100824.

Spelda:2012:AHS

[54] Daniel Špelda. Anachronisms in the history of science: an attempt at a typology. Almagest: International Journal for the History of Scientific Ideas, 3(2):90–119, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100825.

Doomen:2012:UE

[55] Jasper Doomen. Understanding and explaining. Almagest: International Journal for the History of Scientific Ideas, 3(2):120–135, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100826.

Postnikov:2012:BRG

[56] Alexey V. Postnikov. Book review: George Tolias, Mapping Greece, 1420– 1800: a History. Maps in the Margarita Samourkas Collection. Almagest: International Journal for the History of Scientific Ideas, 3 (2):138–142, 2012. CODEN ???? ISSN 1792-2593. URL https:/

/www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100827.

Loureiro:2012:BRN

[57] Rui Manuel Loureiro. Book review: Noël Golvers, Portuguese Books and Their Readers in the Jesuit Mission of China (17th–18th centuries). Almagest: International Journal for the History of Scientific Ideas, 3 (2):143–144, 2012. CODEN ???? ISSN 1792-2593. URL https:/

/www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100828.

Shouguan:2012:BRC

[58] Wang Shouguan. Book review: Chinese Academy of Sciences, A Great Series on the History of Astronomy in China, 10 volumes. Almagest: International Journal for the History of Scientific Ideas, 3(2):145–166, 2012. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100829.

Anonymous:2012:BMGb

[59] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 3(2):168–171, 2012. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.69010603017108027805030007050077.

Anonymous:2013:FMTa

[60] Anonymous. Front matter (‘Title page’, ‘Editorial Board’, ‘table of contents’). Almagest: International Journal for the History of Scientific Ideas, 4(1):1–3, 2013. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.88080103050608820400060100080201.

Malpangotto:2013:CUP

[61] Michela Malpangotto. La critique de l’univers de Peurbach développée par Albert de Brudzewo a-t-elle influencé Copernic ? Un nouveau regard sur les réflexions astronomiques au XVe siècle. (French) [Did the critique of Peurbach’s universe developed by Albert de Brudzewo influence Copernicus? A new look at astronomical reflections in the 15th century]. Almagest: International Journal for the History of Scientific Ideas, 4(1):4–61, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100982.

Golvers:2013:NMR

[62] Noël Golvers. A note on the “machine of Roemer” in late-17th century China, Antoine Thomas, S.J., and the first contacts of Ferdinand Verbiest, S.J., with the Jesuits in Paris. Almagest: International Journal for the History of Scientific Ideas, 4(1):62–73, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100983.

Brentjes:2013:NKI

[63] Sonja Brentjes. Narratives of knowledge in Islamic societies: What do they tell us about scholars and their contexts? Almagest: International Journal for the History of Scientific Ideas, 4(1):74–95, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100984.

Nugayev:2013:PCT

[64] Rinat M. Nugayev. The Ptolemy– Copernicus transition: Intertheoretic context. Almagest: International Journal for the History of Scientific Ideas, 4(1):96–119, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.100985.

Anonymous:2013:BMGa

[65] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 4(1):120–128, 2013.

CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.5555/J.ALMAGEST.
                                        6.88870704040500860303027407088800.

Anonymous:2013:FMTb

[66] Anonymous. Front matter (‘Title page’, ‘Editorial Board’, ‘table of contents’). Almagest: International Journal for the History of Scientific Ideas, 4(2):1–3, 2013. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.5555/J.ALMAGEST.
                                        6.72080602068703050203020304010182.

Wright:2013:AMC

[67] M. T. Wright. The Antikythera mechanism: Compound gear-trains for planetary indications. Almagest: International Journal for the History of Scientific Ideas, 4(2):4–31, 2013. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.1.103717.

Postnikov:2013:RMB

[68] Alexey V. Postnikov. The Russian mission to Bukhara in 1842 and the fate of the English emissaries Charles Stoddart and Arthur Conolly. Almagest: International Journal for the History of Scientific Ideas, 4(2):32–59, 2013. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.1.103718.

Petkovic:2013:ALI

[69] Tomislav Petković. The achievement, legacy, intuition, and cosmopolitanism of Nikola Tesla. Almagest: International Journal for the History of Scientific Ideas, 4(2):60–85, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.1.103719.

Pisano:2013:CMS

[70] Raffaele Pisano and Danilo Capecchi. Conceptual and mathematical structures of mechanical science in the Western Civilization around the 18th century. Almagest: International Journal for the History of Scientific Ideas, 4(2):86–120, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.1.103720.

Argyrakis:2013:EWT

[71] Vaios Argyrakis. The evolution of wind theory and the concept of exhalation in the context of ancient Greek thought. Almagest: International Journal for the History of Scientific Ideas, 4(2):122–136, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.1.103721.

Katsiampoura:2013:BRI

[72] Gianna Katsiampoura. Book review: Ioannis Telelis (ed.) Georgios Pachymeres, Philosophia. Book 5, Commentary in Aristotle’s Meteorologica. [Greek title], editio princeps. Almagest 4(2) 140–141. Almagest: International Journal for the History of Scientific Ideas, 4(2):140–141, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.1.103722.

Vlahakis:2013:BRI

[73] George N. Vlahakis. Book review: Iannis Karas, [Greek title] (Theophilos Kairis, the scientist, the philosopher. A prominent figure of the modern Greek letters). Almagest: International Journal for the History of Scientific Ideas, 4(2):141–143, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.1.103723.

Vlahakis:2013:BRM

[74] George N. Vlahakis. Book review: Manolis Patiniotis, [Greek title] (Elements of Natural Philosophy The Greek scientific thought in the 17th and 18th century). Almagest: International Journal for the History of Scientific Ideas, 4(2):144–146, 2013. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.1.103724.

Anonymous:2013:BMGb

[75] Anonymous. Back matter (’Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 4(2):147–151, 2013. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.87017705020105030506060808020203.

Anonymous:2014:FMTa

[76] Anonymous. Front matter (‘Title page’, ‘Editorial Board’, ‘table of contents’). Almagest: International Journal for the History of Scientific Ideas, 5(1):1–3, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.00050708070607000572060207870603.

Meeus:2014:FEU

[77] Dominique Meeùs. Friedrich Engels and the unveiling of the historical dimension of the physical world: science and dialectics. Almagest: International Journal for the History of Scientific Ideas, 5(1):4–13, 2014. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.102470.

Mazurkiewicz:2014:WMA

[78] Stany Mazurkiewicz. Which materialism? Against which idealism? Dialectical thinking of nature according to Hegel, Engels and Schelling. Almagest: International Journal for the History of Scientific Ideas, 5(1):14–21, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.102469.

Viltanioti:2014:PIR

[79] Irini-Fotini Viltanioti. Plato on iron reduction technique (ti. 60d2-5). Almagest: International Journal for the History of Scientific Ideas, 5(1):22–31, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.102471.

Golvers:2014:FVO

[80] Noël Golvers. Ferdinand Verbiest’s 1668 observation of an unidentified celestial phenomenon in Peking, its lost Chinese description and some parallel observations, especially in Korea. Almagest: International Journal for the History of Scientific Ideas, 5(1):32–51, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.102472.

Koutalis:2014:MDB

[81] Vangelis Koutalis. Making discoveries for a better life vs. bringing fruits to the national treasury: Davy, Babbage, Brewster and the (ongoing) struggle for the soul of science. Almagest: International Journal for the History of Scientific Ideas, 5(1):52–109, 2014. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.102473.

Anonymous:2014:BMGa

[82] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 5(1):110–113, 2014. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.5555/J.ALMAGEST.
                                        6.74087603780777000102060205717804.

Anonymous:2014:FMTb

[83] Anonymous. Front matter (‘Title page’, ‘Editorial Board’, table of contents’). Almagest: International Journal for the History of Scientific Ideas, 5(2):1–3, 2014. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.5555/J.ALMAGEST.
                                        6.75070201030205057504070200020203.

Vlahakis:2014:SCI

[84] George N. Vlahakis. Scientific cosmopolitanism: an introduction. Almagest: International Journal for the History of Scientific Ideas, 5(2):4–7, 2014. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.103563.

Debarbat:2014:JKB

[85] Suzanne Débarbat and Simone Dumont. Johann Karl Buckhardt, un Allemand de Gotha à Paris. (French) [Johann Karl Buckhardt, a German from Gotha in Paris]. Almagest: International Journal for the History of Scientific Ideas, 5(2):8–25, 2014. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.103564.

Meyer-Spasche:2014:COB

[86] Rita Meyer-Spasche. Cosmopolitan Oscar Buneman (1913–1993): his serpentine path from Milan to Stanford. Almagest: International Journal for the History of Scientific Ideas, 5(2):26–39, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.103565.

Wahl:2014:BCN

[87] Charlotte Wahl. Between cosmopolitanism and nationalism. The role of expatriates in the dissemination of Leibniz’s differential calculus. Almagest: International Journal for the History of Scientific Ideas, 5(2):40–68, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.103566.

Muursepp:2014:DNE

[88] Peeter Müürsepp. Dawn of the New Enlightenment. Almagest: International Journal for the History of Scientific Ideas, 5(2):70–83, 2014. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.103567.

Terdimou:2014:UEM

[89] Maria Terdimou. The unsolved equation: Mathematics at the University of Athens during the 19th century. Almagest: International Journal for the History of Scientific Ideas, 5(2):84–101, 2014. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.103568.

Casado:2014:DSP

[90] Carlos M. Madrid Casado. The depiction of science in the paintings of the Museo del Prado. Science and art in the Spanish Empire (16th–18th century). Almagest: International Journal for the History of Scientific Ideas, 5(2):102–124, 2014. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.103569.

Nicolaidis:2014:BRJ

[91] Efthymios Nicolaidis. Book review: Jean-Luc Fournet and Anne Tihon, Conformément aux observations d Hipparque: le papyrus Fouad inv. 267A. (Annex by Raymond Mercier). (French) [Jean-Luc Fournet and Anne Tihon, In accordance with the observations of Hipparchus: the Fouad papyrus inv. 267A]. Almagest: International Journal for the History of Scientific Ideas, 5(2):128–130, 2014. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.103570.

Delli:2014:BRM

[92] Eudoxie Delli. [Book review]: Michel Blay, Dieu, la nature et l’homme. L’originalité de l’Occident. (French)

[Michel Blay, God, nature and man. The originality of the West]. Almagest: International Journal for the History of Scientific Ideas, 5(2):130–139, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.103571.

Katsiampoura:2014:BRM

[93] Gianna Katsiampoura. Book review: Matteo Martelli, The Four Books of Pseudo-Democritus, Society for the History of Alchemy and Chemistry, Maney Publishing. Almagest: International Journal for the History of Scientific Ideas, 5(2):139–141, 2014. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.103572.

Anonymous:2014:BMGb

[94] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 5(2):142–143, 2014. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
6.81010701740204020502060605040005.

Anonymous:2015:FMTa

[95] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 6(1):1–3, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.5555/J.ALMAGEST.4.00001.

Petrovic:2015:BPS

[96] Aleksandar Petrović and Tomislav Petković. Boscovichianism pure and simple. A foreword. Almagest: International Journal for the History of Scientific Ideas, 6(1):4–7, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.107460.

Petrovic:2015:PBT

[97] Aleksandar Petrović. The presence of Boscovich’s Theoria Philosophiae Naturalis and the absence of its translations. Almagest: International Journal for the History of Scientific Ideas, 6(1):8–21, 2015. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.107468.

Stavinschi:2015:BRT

[98] Magda Stavinschi. Boscovich on the Romanian territories. Almagest: International Journal for the History of Scientific Ideas, 6(1):22–31, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.107461.

Savkovic:2015:OPR

[99] Nada Savkovic. Occasional poetry by Roger Boscovich. Almagest: International Journal for the History of Scientific Ideas, 6(1):32–47, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.107462.

Ursic:2015:BDB

[100] Marko Uršič. Boscovich’s distinction between the potential and the actual space from the standpoint of modern search for the union of mind and nature. Almagest: International Journal for the History of Scientific Ideas, 6(1):48–63, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.107463.

Lukan:2015:RBQ

[101] Peter Lukan. Roger Boscovich and the quantum mechanical combination of dynamic and statistical laws. Almagest: International Journal for the History of Scientific Ideas, 6(1):64–78, 2015. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.107464.

Petkovic:2015:RJB

[102] Tomislav Petković. R. J. Boscovich’s achievement in natural philosophy in relation to the development of modern particle physics. Almagest: International Journal for the History of Scientific Ideas, 6(1):80–103, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.107465.

Zorlu:2015:EHL

[103] Tuncay Zorlu. Engineering and humour in the Late Ottoman Empire: The journal SHAQA. Almagest: International Journal for the History of Scientific Ideas, 6(1):104–124, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.107466.

Pourprix:2015:MPA

[104] Bernard Pourprix. La mécanique physique au sens de Joseph Boussinesq. (French) [Physical mechanics in the sense of Joseph Boussinesq]. Almagest: International Journal for the History of Scientific Ideas, 6(1):126–155, 2015. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.107467.

Anonymous:2015:BMGa

[105] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 6(1):156–159, 2015. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.4.00003.

Anonymous:2015:FMTb

[106] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 6(2):1–3, 2015. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.5.109939.

Tihon:2015:RSA

[107] Anne Tihon. Remarques sur les scolies anciennes de l’Almageste. (French) [Remarks on the ancient scholia of the Almagest]. Almagest: International Journal for the History of Scientific Ideas, 6(2):4–41, 2015. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.109660.

Gasca:2015:JFF

[108] Millán Gasca. “Une jeune fille fort sérieuse”: Suzanne Lambin (1902– 2008), a pioneer in biomedicine and biomathematics. Almagest: International Journal for the History of Scientific Ideas, 6(2):42–68, 2015. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.109661.

Vandersmissen:2015:DEF

[109] Jan Vandersmissen. Débats européens sur les formes de la vie marine au Siècle des Lumières. (French) [European debates on the forms of marine life during the Age of Enlightenment]. Almagest: International Journal for the History of Scientific Ideas, 6(2):70–83, 2015. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.109662.

Sigrist:2015:CTF

[110] René Sigrist and Alexander Moutchnik.

Entre Ciel et Terre : les fonctions de l’astronomie dans la Russie du 18e siècle. (French) [Between Heaven and Earth: the functions of astronomy in 18th century Russia]. Almagest: International Journal for the History of Scientific Ideas, 6(2):84–125, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.109663.

Pisano:2015:FAS

[111] Raffaele Pisano and Paolo Bussotti.

Fibonacci and the Abacus Schools in Italy. Mathematical conceptual streams — education and its changing relationship with society. Almagest: International Journal for the History of Scientific Ideas, 6(2):126–164, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.109664.

Danova:2015:BRA

[112] Nadia Danova. Book review: Alexandre Kostov, [Cyrillic title] (From Craft to Profession. Engineering in the Balkans from the Beginning of the 18th Century to World War I), Sofia, Paradigma, 2015, 412 p. Almagest: International Journal for the History of Scientific Ideas, 6(2):169–173, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.109667.

Anonymous:2015:BMGb

[113] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 6(2):174–176, 2015. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.5555/J.ALMAGEST.5.109940.

Anonymous:2016:FMC

[114] Anonymous. Front matter (’Contents’). Almagest: International Journal for the History of Scientific Ideas, 7(1):1–3, 2016. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.5555/J.ALMAGEST.
5.111618.

Anonymous:2016:GPI

[115] Anonymous. General preface to the inscriptions. Almagest: International Journal for the History of Scientific Ideas, 7(1):4–35, 2016. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.110732.

Jones:2016:HBG

[116] A. Jones. Historical background and general observations. Almagest: International Journal for the History of Scientific Ideas, 7(1):36–67, 2016. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.110733.

Bitsakis:2016:FDP

[117] Y. Bitsakis and A. Jones. The front dial and Parapegma inscriptions. Almagest: International Journal for the History of Scientific Ideas, 7(1):68–137, 2016. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.110734.

Anastasiou:2016:BDB

[118] M. Anastasiou, Y. Bitsakis, A. Jones, J. M. Steele, and M. Zafeiropoulou. The back dial and back plate inscriptions. Almagest: International Journal for the History of Scientific Ideas, 7(1):138–215, 2016. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.110735.

Bitsakis:2016:BCI

[119] Y. Bitsakis and A. Jones. The back cover inscription. Almagest: International Journal for the History of Scientific Ideas, 7(1):216–249, 2016. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.110736.

Anastasiou:2016:FCI

[120] A. Anastasiou, Y. Bitsakis, A. Jones, X. Moussas, A. Tselikas, and M. Zafeiropoulou. The front cover inscription. Almagest: International Journal for the History of Scientific Ideas, 7(1):250–297, 2016. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.110737.

Anonymous:2016:SI

[121] Anonymous. Supplementary illustrations. Almagest: International Journal for the History of Scientific Ideas, 7(1):299–310, 2016. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.111619.

Anonymous:2016:BMGa

[122] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 7(1):312–315, 2016. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.5.111621.

Anonymous:2016:FMT

[123] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 7(2):C1, 2016. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.4.2017001.

Meyer-Spasche:2016:PPO

[124] Rita Meyer-Spasche and Rolf Tomas Nossum. Persecution and patronage: Oscar Buneman’s years in Britain. Almagest: International Journal for the History of Scientific Ideas, 7(2):4–21, 2016. ISSN

                                        1792-2593.  URL https://pure.
                                        mpg.de/rest/items/item_2403602/
                                        component/file_2403630/content;
                                         https://www.brepolsonline.net/
                                        doi/10.1484/J.ALMAGEST.5.112684.

Pechenkin:2016:SB

[125] Alexander Pechenkin. The story of a book. Almagest: International Journal for the History of Scientific Ideas, 7(2):23–48, 2016. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.112685.

Mukhin:2016:DDC

[126] Ravil Mukhin. The development of dynamics in the 20th century. Kolmogorov–Arnold–Moser (KAM) theory. Almagest: International Journal for the History of Scientific Ideas, 7(2):51– 72, 2016. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.112686.

Bonnin:2016:PDC

[127] Jérôme Bonnin, Jean-Marie Michel, and Jean-Yves Thiant. Le pelecinum de Doumet sur la commune de Châteauvert (Var). (French) [The Doumet pelecinum in the commune of Châteauvert (Var)]. Almagest: International Journal for the History of Scientific Ideas, 7(2):75–103, 2016. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.112687.

Gapochenko:2016:APH

[128] Svetlana Gapochenko and Halina Khlyap. Aesthetic principles in the history of physics. Almagest: International Journal for the History of Scientific Ideas, 7(2):105–119, 2016. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.112688.

Anonymous:2016:BMGb

[129] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 7(2):120–C3, 2016. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.4.2017002.

Anonymous:2017:FMTa

[130] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 8(1):c1, 2017. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.4.2017005.

Kolchinsky:2017:NVR

[131] Eduard I. Kolchinsky. Nikola I. Vavilov in the realm of historical and scientific discussions. Almagest: International Journal for the History of Scientific Ideas, 8(1):5–36, 2017. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.113696.

Sigrist:2017:FSD

[132] René Sigrist and Alexander Moutchnik. Les fondements sociaux du premier essor de la botanique en Russie, 1700— 1830. (French) [The social foundations of the first development of botany in Russia, 1700—1830]. Almagest: International Journal for the History of Scientific Ideas, 8(1):39–75, 2017. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.113697.

Golvers:2017:CKB

[133] Noël Golvers. Circulation of knowledge between Europe and China in the 17th–18th centuries: The role of material objects, from gadgets to instruments. Almagest: International Journal for the History of Scientific Ideas, 8(1):77–117, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.113698.

Carman:2017:HPS

[134] Christián C. Carman. Heraclides of Pontus on the solar anomaly. Almagest: International Journal for the History of Scientific Ideas, 8(1):119–125, 2017. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.113699.

Kilakos:2017:BRJ

[135] Dimitris Kilakos. Book review: Jimena Canales, The Physicist and the Philosopher: Einstein, Bergson and the Debate that Changed our Understanding of Time. Almagest: International Journal for the History of Scientific Ideas, 8(1):129–132, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.113700.

Nicolaidis:2017:BRA

[136] Efthymios Nicolaidis. Book review:

Alexander Jones, A portable Cosmos. Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World. Almagest: International Journal for the History of Scientific Ideas, 8(1):133–137, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.113701.

Anonymous:2017:BME

[137] Anonymous. Back matter (’Erratum’, ’Le pelecinum de Doumet sur la commune de Châteauvert (Var) (Almagest 7.2, pp. 75–103)’, ‘guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 8(1):139–142, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.5555/J.ALMAGEST.5.114141.

Anonymous:2017:FMTb

[138] Anonymous. Front matter (‘Table of contents’, ‘foreword’). Almagest: International Journal for the History of Scientific Ideas, 8(2):C1, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.5555/J.ALMAGEST.4.2017008.

Brooke:2017:AWE

[139] John Hedley Brooke and Ronald L. Numbers. Into all the world: expanding the history of science and religion beyond the Abrahamic faiths. Almagest: International Journal for the History of Scientific Ideas, 8(2):9–26, 2017. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.114929.

Caudano:2017:CAF

[140] Anne-Laurence Caudano. Cosmography, asceticism and female patronage in Late Byzantine and Slavic miscellanies. Almagest: International Journal for the History of Scientific Ideas, 8(2):29–47, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.114930.

Shank:2017:APA

[141] Michael H. Shank. The Almagest, politics, and apocalypticism in the conflict between George of Trebizond and Cardinal Bessarion. Almagest: International Journal for the History of Scientific Ideas, 8(2):49–83, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.114931.

Vafea:2017:AIS

[142] Flora Vafea. The astronomical instruments in Saint Catherine’s iconography at the Holy Monastery of Sinai. Almagest: International Journal for the History of Scientific Ideas, 8(2):85–109, 2017. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.114932.

Petkovic:2017:RJB

[143] Tomislav Petković. R. J. Bošković as European scientist and theologian at work on the bridges between science and religion. Essay on Bošković’s top concept and epistemological interpretation of God. Almagest: International Journal for the History of Scientific Ideas, 8(2):111–134, 2017. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.114933.

Tampakis:2017:OSG

[144] Kostas Tampakis. Orthodoxy and science in the Greek State, 1830–1939. Almagest: International Journal for the History of Scientific Ideas, 8(2):137–157, 2017. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.114934.

Anonymous:2017:BMG

[145] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 8(2):158–160, 2017. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.4.2017009.

Anonymous:2018:FMTa

[146] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 9(1):C1, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.4.2018001.

Golvers:2018:MIC

[147] Noël Golvers. Mathematical instruction at the Colégio das Artes in Coimbra: Adam Aigenler’s “rota astronomica” (1672). Almagest: International Journal for the History of Scientific Ideas, 9(1):5–16, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.116016.

Sheynin:2018:SFB

[148] Oscar Sheynin. ‘Saint Fedos’. a biography of Feodosy Nikolaevich Krasovsky (1878–1948). Almagest: International Journal for the History of Scientific Ideas, 9(1):19–30, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.116017.

Prapas:2018:ECS

[149] Fotios Prapas. Elements from the cosmology of Saint Maximus the Confessor. Man as mediator of salvation of Cosmos. Almagest: International Journal for the History of Scientific Ideas, 9(1):33–49, 2018. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.116018.

Andriotis:2018:TGT

[150] Antonios N. Andriotis. The theory of “[Greek title]” (Pantahekineton) of Benjamin Lesvios. Almagest: International Journal for the History of Scientific Ideas, 9(1):51–67, 2018. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.116019.

Wang:2018:SGS

[151] Kai Wang. Scientific gentry and socialisation of Western Science in China’s modernisation during ’Selfstrengthening’ Movement (1860–1895). Almagest: International Journal for the History of Scientific Ideas, 9(1):69–95, 2018. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.116020.

Zaytsev:2018:POM

[152] Evgeny Zaytsev. The priority of object over method in Early Greek mathematics. Almagest: International Journal for the History of Scientific Ideas, 9(1):97–119, 2018. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.116021.

Anonymous:2018:BMGa

[153] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 9(1):120–123, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.5555/J.ALMAGEST.4.2018002.

Anonymous:2018:FMTb

[154] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 9(2):C1–C3, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.4.2019001.

Brooke:2018:OSV

[155] John Hedley Brooke. Orthodoxy and science: Varieties of engagement and indifference. Almagest: International Journal for the History of Scientific Ideas, 9(2):4–11, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.116767.

Kilakos:2018:RIC

[156] Dimitris Kilakos. Religion and ideological confrontations in early Soviet mathematics: The case of P. A. Nekrasov. Almagest: International Journal for the History of Scientific Ideas, 9(2):12–39, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.116768.

Howell:2018:RSA

[157] Christopher Howell. The Rose and the Stag. An American Orthodox conversation on modernity, science, and biblical interpretation. Almagest: International Journal for the History of Scientific Ideas, 9(2):40–59, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.116769.

Tihon:2018:DHP

[158] Anne Tihon. Le diagramme des horizons et les prosneuses des éclipses dans l’astronomie de Ptolémée. (French) [The diagram of the horizons and the proneurs of eclipses in Ptolemy’s astronomy]. Almagest: International Journal for the History of Scientific Ideas, 9(2):60–87, 2018. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.116770.

Golvers:2018:NJT

[159] Noël Golvers. New Jesuit testimonies from the Far East on the comet of March 1668: the diary of Visitor Luis da Gama (Macau) and a letter of Ferdinand Verbiest (Peking). Almagest: International Journal for the History of Scientific Ideas, 9(2):88–99, 2018. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.116771.

Vasilescu:2018:HPV

[160] Andrei Vasilescu and Magda Stavinschi. 1877 — Henri Poincaré : le voyage d’étude et la fabrication de l’acier dans le Banat. (French) [1877 — Henri Poincaré: the study trip and steel manufacturing in Banat]. Almagest: International Journal for the History of Scientific Ideas, 9(2):100–117, 2018. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.116772.

Zorlu:2018:BRE

[161] Tuncay Zorlu. Book review: Ekmeleddin İhsanoğlu, Osmanlı Bilim Mirası (Ottoman Scientific Heritage). Almagest: International Journal for the History of Scientific Ideas, 9(2):120–124, 2018. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.116773.

Daniele:2018:BRC

[162] Stefano Daniele. Book review: C. Guerra, Lavoisier e Parthenope. Contributo ad una storia della chimica del Regno di Napoli (Italian) [C. Guerra, Lavoisier and Parthenope. Contribution to a history of chemistry in the Kingdom of Naples]. Almagest: International Journal for the History of Scientific Ideas, 9(2):124–127, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.116774.

Katsiampoura:2018:BRK

[163] Gianna Katsiampoura. Book review: Kostas V. Krimbas, Darwinism in Greece. Almagest: International Journal for the History of Scientific Ideas, 9(2):127–129, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.116775.

Anonymous:2018:BMGb

[164] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 9(2):130–133, 2018. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.4.2019002.

Anonymous:2019:FMTa

[165] Anonymous. Front matter (‘Table of contents’, ‘Efthymios Nicolaidis and Constantine Skordoulis’, ‘ten years of Almagest’). Almagest: International Journal for the History of Scientific Ideas, 10(1):1–5, 2019. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.4.2019005.

Vlahakis:2019:SII

[166] George N. Vlahakis. Science and imagination: an introduction. Almagest: International Journal for the History of Scientific Ideas, 10(1):6–9, 2019. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.118328.

Hellberg:2019:CET

[167] Dustin Hellberg. Close encounters of the tertiary kind: Science fiction as tertiary epic. Almagest: International Journal for the History of Scientific Ideas, 10(1):10–23, 2019. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.118329.

Altashina:2019:PEP

[168] Veronika Altashina. Probability and expectation in Pascal’s Pensées. Almagest: International Journal for the History of Scientific Ideas, 10(1):24–36, 2019. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.118330.

Godel:2019:WSR

[169] Rainer Godel. Why Samuel Richardson’s Clarissa became a role model. On Johann Georg Zimmermann’s biography of Albrecht von Haller. Almagest: International Journal for the History of Scientific Ideas, 10(1):38–50, 2019. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.118331.

Berg:2019:EVM

[170] Gunhild Berg. Experimenters versus magicians: Poetic strategies of an intertextual rivalry in 19th century German fiction. Almagest: International Journal for the History of Scientific Ideas, 10(1):52–65, 2019. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.118332.

Denissi:2019:GTT

[171] Sophia Denissi. Gulliver’s ‘travels’ through the European sea to the land of 19th century Greek translations. Almagest: International Journal for the History of Scientific Ideas, 10(1):66–77, 2019. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.118333.

Couleau:2019:SFI

[172] Christèle Couleau. Is science a foil? Indirect powers of the novelistic speech in Jules Verne’s work. Almagest: International Journal for the History of Scientific Ideas, 10(1):78–90, 2019. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.118334.

Lazos:2019:AMF

[173] Panagiotis Lazos. Around the Moon. False scientific ideas in the famous book of Jules Verne. Almagest: International Journal for the History of Scientific Ideas, 10(1):92–105, 2019. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.118335.

Vlahakis:2019:PIV

[174] George N. Vlahakis. Political and ideological views of Jules Verne: a consideration through his books. Almagest: International Journal for the History of Scientific Ideas, 10(1):106–115, 2019. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.118336.

Terdimou:2019:EMA

[175] Maria Terdimou. Elements of mathematics and astronomy in the work of Jules Verne. Almagest: International Journal for the History of Scientific Ideas, 10(1):116–130, 2019. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.118337.

Anonymous:2019:BME

[176] Anonymous. Back matter (‘Errata’, ‘guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 10(1):132–135, 2019. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
4.2019006.

Anonymous:2019:FMTb

[177] Anonymous. Front matter (“Ten years of Almagest”, “Table of contents”). Almagest: International Journal for the History of Scientific Ideas, 10(2):1–7, 2019. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.120041.

Sigrist:2019:FGR

[178] René Sigrist. The first generations of Russian chemists (1700–1830). Between pharmacy, technical know-how and academia. Almagest: International Journal for the History of Scientific Ideas, 10(2):8–51, 2019. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.119493.

Jimenez:2019:ASD

[179] Carlos Alvarez Jiménez, Vincent Jullien, and Carmen Martínez-Adame. On the analytic and synthetic demonstrations in Fermat’s work on the law of refraction. Almagest: International Journal for the History of Scientific Ideas, 10(2):52–85, 2019. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.119494.

Kharlamova:2019:ABP

[180] Vera I. Kharlamova. About bibliographical projects during the formation of the mathematical community in Europe. Portuguese mathematicians in international bibliography at the turn of the XIX–XX centuries. Almagest: International Journal for the History of Scientific Ideas, 10(2):86–98, 2019. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.119495.

Kilakos:2019:HBC

[181] Dimitris Kilakos. Higgs boson and the Cosmos: a philosophical reappraisal of the authoritative Catholic and Greek– Orthodox perspectives. Almagest: International Journal for the History of Scientific Ideas, 10(2):100–121, 2019.

CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.119496.

Cerna:2019:BIS

[182] Jana Černá. Beyond the “Irrational and superstitious”: The New World, Spanish “Hunters of nature’s secrets”, and transformations of science in the Early Modern Era. Almagest: International Journal for the History of Scientific Ideas, 10(2):122–145, 2019. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.119497.

Anonymous:2019:BMG

[183] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 10(2):146–153, 2019. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120042.

Anonymous:2020:FMTa

[184] Anonymous. Front matter (“Table of contents”). Almagest: International Journal for the History of Scientific Ideas, 11(1):1–5, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.121327.

Delli:2020:F

[185] Eudoxie Delli and Efthymios Nicolaidis. Foreword. Almagest: International Journal for the History of Scientific Ideas, 11(1):6–11, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120850.

Delli:2020:I

[186] Eudoxie Delli and Efthymios Nicolaidis. Introduction. Almagest: International Journal for the History of Scientific Ideas, 11(1):12–16, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120851.

Meunier:2020:PPB

[187] Florence Meunier. De Platon et Philon à Basile de Césarée et Georges Pisidès: les métamorphoses du démiurge dans son œuvre devenue outil de perfectionnement chrétien. (French) [From Plato and Philo to Basil of Caesarea and George Pisides: the metamorphoses of the demiurge in his work which has become a tool for Christian improvement]. Almagest: International Journal for the History of Scientific Ideas, 11(1):17–41, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120852.

Nicolaidis:2020:PHO

[188] Efthymios Nicolaidis. Platon et Hipparque ont lu la Bible: l’Hexaéméron de Jean Philopon. (French) [Plato and Hipparchus read the Bible: Hexaemeron by John Philopon]. Almagest: International Journal for the History of Scientific Ideas, 11(1):42–55, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120853.

Delli:2020:STL

[189] Eudoxie Delli. Survie et transformations de la littérature hexaémérique. Entre littérature, théologie et philosophie : Le cas de Georges de Pisidie. (French)

[Survival and transformations of Hexaamerican literature. Between literature, theology and philosophy: The case of George of Pisidia]. Almagest: International Journal for the History of Scientific Ideas, 11(1):56–79, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120854.

Lamy:2020:MCQ

[190] Alice Lamy. La matière de la création dans quatre commentaires Carolingiens sur la Genèse (I, 1-14) : une origine du monde entre forme et ténèbres. (French) [The matter of creation in four Carolingian commentaries on Genesis (I, 1-14): an origin of the world between form and darkness]. Almagest: International Journal for the History of Scientific Ideas, 11(1):80–106, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120855.

Brouwer:2020:RGL

[191] Christian Brouwer. Robert Grosseteste: la lumière et les fondements du monde. (French) [Robert Grosseteste: the light and the foundations of the world]. Almagest: International Journal for the History of Scientific Ideas, 11(1):108–119, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.120856.

Delhove:2020:CQF

[192] Arnaud Delhove. De ce qui fut créé le sixième jour … Discours sur l’homme dans le Candélabre des Sanctuaires de Bar Hebræus (II, 3, 6). (French) [Of what was created on the sixth day … Discourse on man in the Candelabra of the Sanctuaries of Bar Hebræus (II, 3, 6)]. Almagest: International Journal for the History of Scientific Ideas, 11(1):120–136, 2020. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.120857.

Delli:2020:BSS

[193] Eudoxie Delli. Bibliographie secondaire sélective sur les Hexaéméra et les thématiques rattachées. (French) [Selective secondary bibliography on the Hexaemera and related themes]. Almagest: International Journal for the History of Scientific Ideas, 11(1):138–156, 2020. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.120858.

Anonymous:2020:BMA

[194] Anonymous. Back matter (“Annexe: Le texte de Genèse 1,1-2,4 [traduction française]”, “Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 11(1):158–167, 2020. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.121328.

Anonymous:2020:FMTb

[195] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 11(2):1–3, 2020. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.124459.

Savoie:2020:EGA

[196] Denis Savoie. Étude gnomonique de l’anneau astronomique de Philippes. (French) [Gnomonic study of the astronomical ring of Philippi]. Almagest: International Journal for the History of Scientific Ideas, 11(2):4–22, 2020. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.123108.

Dalong:2020:SLT

[197] Lu Dalong. The solar and lunar theories of Qizheng Tuibu (1477). Almagest: International Journal for the History of Scientific Ideas, 11(2):24–56, 2020. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.123109.

Kubbinga:2020:NMD

[198] Henk Kubbinga. Nicolaus Mulerius (1564–1630): Dissimulating Copernicanism in the revolting Dutch Republic. Almagest: International Journal for the History of Scientific Ideas, 11(2):58–83, 2020. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.123110.

Bonfim:2020:HMS

[199] Sabrina Helena Bonfim and Sergio Roberto Nobre. Historical mathematical study about vector calculus introduction in Brazil: First notes. Almagest: International Journal for the History of Scientific Ideas, 11(2):84–110, 2020. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.

net/doi/epdf/10.1484/J.ALMAGEST.
5.123111.

Sfoini:2020:TSA

[200] Alexandra Sfoini. Traduire les sciences au cours des Lumières néohélleniques : questions de langue et de terminologie (1750–1832). (French) [Translating the sciences during the Neohellenic Enlightenment: questions of language and terminology (1750–1832)]. Almagest: International Journal for the History of Scientific Ideas, 11(2):112–123, 2020. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.123112.

Vlahakis:2020:BRE

[201] George N. Vlahakis. Book review: Ekmeleddin İhsanoğlu: The House of Sciences. The First Modern University in the Muslim World. Almagest: International Journal for the History of Scientific Ideas, 11(2):126–127, 2020. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.123113.

Anonymous:2020:BMG

[202] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 11(2):128–131, 2020. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.124460.

Anonymous:2021:FMT

[203] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 12(??):1–3, 2021. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.128362.

Park:2021:I

[204] Hyung Wook Park and Ronald L. Numbers. Introduction. Almagest: International Journal for the History of Scientific Ideas, 12(??):4–11, 2021. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.125383.

Park:2021:PCS

[205] Hyung Wook Park. Practicing Creationism: Science and the new religious practices in South Korea. Almagest: International Journal for the History of Scientific Ideas, 12(??):12–39, 2021. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.125384.

Edis:2021:TMI

[206] Taner Edis. The Turkish model of Islam Creationism. Almagest: International Journal for the History of Scientific Ideas, 12(??):40–65, 2021. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.125385.

Godart:2021:ETS

[207] G. Clinton Godart. ‘Evolutionary theory is the superstition of modernity’: Antievolutionary thought in wartime Japan. Almagest: International Journal for the History of Scientific Ideas, 12(??):66–91, 2021. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.125386.

Stenhouse:2021:CEN

[208] John Stenhouse. Creationists and evolutionists in New Zealand, 1800– 2010: Science, religion, politics, and race. Almagest: International Journal for the History of Scientific Ideas, 12(??):92–122, 2021. CODEN ???? ISSN 1792-2593. URL https:/

                                        /www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.125387.

Aechtner:2021:CAA

[209] Thomas Aechtner. Creationism with an Australian accent: Politics, schools, and global exportation. Almagest: International Journal for the History of Scientific Ideas, 12(??):124–148, 2021. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.125388.

Brown:2021:HC

[210] C. Mackenzie Brown. Hindu Creationism. Almagest: International Journal for the History of Scientific Ideas, 12(??):150–191, 2021. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.125389.

Zhang:2021:CEC

[211] Zeng yi Zhang. Creation–evolution controversies in China: a study of intelligent design in social media. Almagest: International Journal for the History of Scientific Ideas, 12(??):192–207, 2021. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.125390.

Nicolaidis:2021:CTO

[212] Efthymios Nicolaidis. Creationism in today’s Orthodox community. Almagest: International Journal for the History of Scientific Ideas, 12(??):208–226, 2021. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.125391.

Nicolaidis:2021:CBK

[213] Efthymios Nicolaidis. Costas B. Krimbas 1932–2021. Almagest: International Journal for the History of Scientific Ideas, 12(??):228–233, 2021. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.126101.

Anonymous:2021:BMG

[214] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 12(??):235–239, 2021. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.128363.

Anonymous:2022:FMTa

[215] Anonymous. Front matter (‘Table of contents’). Almagest: International Journal for the History of Scientific Ideas, 13(1):1–3, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.132258.

Dalong:2022:PLT

[216] Lu Dalong and Han Dongyang. The planetary longitude theory of Qizheng Tuibu (1477). Almagest: International Journal for the History of Scientific Ideas, 13(1):4–41, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.131487.

Carman:2022:GMC

[217] Christián C. Carman. The Great Martian Catastrophe and how Tycho (re- )fixed it. Almagest: International Journal for the History of Scientific Ideas, 13(1):42–57, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.131488.

Chassefiere:2022:LBE

[218] Eric Chassefière. L’aurore boréale enjeu du mécanisme cartésien et la dispute entre Paris et Montpellier, le choix français. (French) [The Northern Lights, the issue of the Cartesian mechanism and the dispute between Paris and Montpellier, the French choice]. Almagest: International Journal for the History of Scientific Ideas, 13(1):58–78, 2022. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.131489.

Crippa:2022:TEM

[219] Davide Crippa. Teaching elementary mathematics at the University of Prague: a study of Latin compendia from the second half of the 18th century. Almagest: International Journal for the History of Scientific Ideas, 13(1):80–141, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.131490.

Anonymous:2022:BR

[220] Anonymous. Book reviews. Almagest: International Journal for the History of Scientific Ideas, 13(1):142–148, 2022. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.131491.

Anonymous:2022:BMGa

[221] Anonymous. Back matter (‘Guidelines for authors’). Almagest: International Journal for the History of Scientific Ideas, 13(1):149–152, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.132259.

Anonymous:2022:FMTb

[222] Anonymous. Front matter (“Table of contents”). Almagest: International Journal for the History of Scientific Ideas, 13(2):1–3, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.133877.

Morfouli:2022:I

[223] Meropi Morfouli. Introduction. Almagest: International Journal for the History of Scientific Ideas, 13(2):4–10, 2022. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.132473.

Jones:2022:PTO

[224] Alexander Jones. Precision of time observation in Greco–Roman astrology and astronomy. Almagest: International Journal for the History of Scientific Ideas, 13(2):12–34, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.132474.

Morfouli:2022:PMD

[225] Meropi Morfouli. La precision dans la mesure du temps et la Nouvelle Philosophie: Étude de cas (Alexandre Koyré & Richard Westfall). (French) [Precision in time measurement and the New Philosophy: Case study (Alexandre Koyré & Richard Westfall)]. Almagest: International Journal for the History of Scientific Ideas, 13(2):36–66, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.132475.

Guicciardini:2022:NAT

[226] N. Guicciardini. Newtonian absolute time vs fluxional time. Almagest: International Journal for the History of Scientific Ideas, 13(2):68–83, 2022. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.132476.

Provost:2022:THI

[227] Jean-Pierre Provost. Three historical issues dealing with time and inertia in relation to present physics. Almagest: International Journal for the History of Scientific Ideas, 13(2):84–94, 2022. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.132477.

Bracco:2022:FOR

[228] Christian Bracco and Jean-Pierre Provost. First-order relativity: From Fresnel’s 1818 “drag” coefficient to Lorentz’s 1895 “local time” followed by Poincaré’s 1900 interpretation, and finally up to Einstein’s 1911–1912 spatially flat metric. Almagest: International Journal for the History of Scientific Ideas, 13(2):96–111, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.132478.

Beisbart:2022:CCR

[229] Claus Beisbart. Can cosmologists really measure the age of the Universe? Almagest: International Journal for the History of Scientific Ideas, 13(2):112–139, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.132479.

Lachieze-Rey:2022:TMT

[230] Marc Lachièze-Rey. Time measurements without time. Almagest: International Journal for the History of Scientific Ideas, 13(2):140–153, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.132480.

Anonymous:2022:BMGb

[231] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 13(2):154–157, 2022. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.133878.

Anonymous:2023:FMTa

[232] Anonymous. Front matter (’Table of Contents’). Almagest: International Journal for the History of Scientific Ideas, 14(1):1–3, January 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.135603.

Voulgaris:2023:ICD

[233] Aristeidis Voulgaris, Christophoros Mouratidis, and Andreas Vossinakis. The initial calibration date of the Antikythera mechanism after the Saros spiral mechanical apokatastasis. Almagest: International Journal for the History of Scientific Ideas, 14(1):4–39, January 2023. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.134609.

Dongyang:2023:PLT

[234] Han Dongyang and Lu Dalong. The planetary latitude theory of Qizheng Tuibu (1477). Almagest: International Journal for the History of Scientific Ideas, 14(1):40–68, January 2023. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.134610.

Ihsanoglu:2023:INA

[235] Ekmeleddin İhsanoğlu. Introductory notes on addressing issues and problems of multireligious-ethnic scholars in the Ottoman science. Almagest: International Journal for the History of Scientific Ideas, 14(1):70–86, January 2023. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.134611.

Schmit:2023:MSO

[236] Christophe Schmit and Éric Chassefière. Météorologie et sciences de l’observatoire à l’Observatoire royal de Paris (1671–1719). (French) [Meteorology and observatory sciences at the Royal Observatory of Paris (1671– 1719)]. Almagest: International Journal for the History of Scientific Ideas, 14(1):88–147, January 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.134612.

Chassefiere:2023:IVB

[237] Éric Chassefière. Interprétation des variations barométriques au tournant du XVIIIe siècle : allègement de l atmosphère par les vapeurs et la pluie. (French) [Interpretation of barometric variations at the turn of the 18th century: lightening of the atmosphere by vapors and rain]. Almagest: International Journal for the History of Scientific Ideas, 14(1):148–173, January 2023. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.
net/doi/epdf/10.1484/J.ALMAGEST.
5.134613.

Magrone:2023:SMJ

[238] Paola Magrone, Ana Millán Gasca, and Ilaria Zannoni. Science for the multitude: Jules Dalsème’s (1845–1904) “Natural geometry” for the education of all. Almagest: International Journal for the History of Scientific Ideas, 14(1):174–209, January 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.134614.

Khaleb:2023:BRE

[239] Dana Khaleb. Book review: Ekmeleddin İhsanoğlu, The Abbasid House of Wisdom between Myth and Reality, New York: Routledge. ISBN: 978-1-032-34745-5 (hardback). Almagest: International Journal for the History of Scientific Ideas, 14(1):210–214, January 2023. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.134615.

Anonymous:2023:BMGa

[240] Anonymous. Back matter (’Guidelines for Authors’). Almagest: International Journal for the History of Scientific Ideas, 14(1):217–221, January 2023. CODEN ???? ISSN 1792-2593.

                                        URL https://www.brepolsonline.
                                        net/doi/epdf/10.1484/J.ALMAGEST.
                                        5.135604.

Anonymous:2023:FMC

[241] Anonymous. Front matter (“Contents”). Almagest: International Journal for the History of Scientific Ideas, 14(2):3–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.137946.

Jullien:2023:F

[242] Vincent Jullien and Carmen Martínez-Adame. Foreward [sic]. Almagest: International Journal for the History of Scientific Ideas, 14(2):6– ??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:/

                                        /www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.136753.

Angleraux:2023:MBM

[243] Caroline Angleraux. De la métaphysique à la biologie, métamorphose conceptuelle des monades. (French) [From metaphysics to biology, conceptual metamorphosis of monads]. Almagest: International Journal for the History of Scientific Ideas, 14

(2):18–??, ????  2023.  CODEN

???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136754.

Bertoldi:2023:LLE

[244] Nicola Bertoldi. L’empirisme logique est-il toujours d’actualité en philosophie de la biologie ?. (French) [Is logical empiricism still relevant in philosophy of biology?]. Almagest: International Journal for the History of Scientific Ideas, 14(2):28–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136755.

Contreras:2023:TMH

[245] Jorge Martínez Contreras. Two moments in the history of primatology: Hanno’s Periplus and Satsue Mito’s legacy. Almagest: International Journal for the History of Scientific Ideas, 14(2):36–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136756.

Ramirez:2023:CEC

[246] Dairon Alfonso Rodríguez Ramírez.

Continuity and exceptionalism in Charles Darwin. Almagest: International Journal for the History of Scientific Ideas, 14(2):46– ??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:/

/www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136757.

Caso:2023:RWH

[247] Juan Manuel Rodríguez Caso and Erica Torrens Rojas. A. R. Wallace and his “evolutionary diagrams”: first Darwinian schemes? Almagest: International Journal for the History of Scientific Ideas, 14 (2):54–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136758.

Valadez:2023:HPC

[248] Bernardo Yáñez Macías Valadez. A historiographical perspective on contemporary biological anthropology. Almagest: International Journal for the History of Scientific Ideas, 14 (2):64–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136759.

Alvarez:2023:SRC

[249] Carlos Alvarez. Some remarks concerning the trisection of an angle. Euclid between Pappus and Viète. Almagest: International Journal for the History of Scientific Ideas, 14 (2):76–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136760.

Dhombres:2023:FHO

[250] Jean Dhombres. Faire l’histoire d’un objet ou celle d’un concept mathématique avec les fonctions mathématiques comme cas d’étude privilégié. (French) [Telling the history of an object or that of a mathematical concept with mathematical functions as a preferred case study]. Almagest: International Journal for the History of Scientific Ideas, 14(2):98–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136761.

Fernandez:2023:IIF

[251] Begoña Fernández. On Itô’s integral and the fundamental theorem of stochastic calculus. Almagest: International Journal for the History of Scientific Ideas, 14(2):136–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136762.

deCastro:2023:PIF

[252] Max Fernández de Castro. Poincaré and impredicativity. Almagest: International Journal for the History of Scientific Ideas, 14(2):150– ??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:/

/www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136763.

Haffner:2023:RRT

[253] Emmylou Haffner. Relectures et réécritures des textes de B. Riemann dans leur édition par R. Dedekind et H. Weber. (French) [Rereadings and rewritings of the texts of B. Riemann in their edition by R. Dedekind and H. Weber]. Almagest: International Journal for the History of Scientific Ideas, 14(2):164–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136764.

auregui-Hernandez:2023:AGS

[254] Anabel áuregui Hernández. On Arnauld’s geometry of straight lines. Almagest: International Journal for the History of Scientific Ideas, 14 (2):176–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136765.

Jullien:2023:DLS

[255] Vincent Jullien. La Dynamica de Leibniz et la science physico-mathématique. (French) [The Dynamica of Leibniz and physico-mathematical science]. Almagest: International Journal for the History of Scientific Ideas, 14 (2):188–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.136766.

Martinez-Adame:2023:SDB

[256] Carmen Martínez-Adame. Similarities and differences between some of the integrals of the 20th century. Almagest: International Journal for the History of Scientific Ideas, 14 (2):210–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.136767.

Perea:2023:CCC

[257] Favio Ezequiel Miranda Perea and Lourdes del Carmen González Huesca. On conceptual changes in computer assisted proofs. Almagest: International Journal for the History of Scientific Ideas, 14(2):222–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.136768.

Rey:2023:CSE

[258] Olivier Rey. Comment la statistique est-elle entrée en physique ?. (French) [How did statistics enter physics?]. Almagest: International Journal for the History of Scientific Ideas, 14 (2):246–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136770.

Waszek:2023:RND

[259] David Waszek. Le rôle des notations dans la découverte de l’analogie des puissances et des différences de Leibniz. (French) [The role of notations in the discovery of the analogy of powers and differences of Leibniz]. Almagest: International Journal for the History of Scientific Ideas, 14(2):256–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136771.

Assada:2023:ICS

[260] Anne-Lise Assada. Incommensurabilité dans le changement scientifique, incommensurabilité des épistémologies réalistes et anti-réalistes. (French) [Incommensurability in scientific change, incommensurability of realist and antirealist epistemologies]. Almagest: International Journal for the History of Scientific Ideas, 14(2):270– ??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:/

/www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136772.

Barahona:2023:EBN

[261] Ana Barahona. Erasing borders: a new look at the historiography of science. Almagest: International Journal for the History of Scientific Ideas, 14(2):280–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.net/doi/epdf/
10.1484/J.ALMA.5.136773.

Wagner:2023:SCS

[262] Pierre Wagner. Sortes de changements scientifiques : changements dans les sciences, changement de la science. (French) [Kinds of scientific changes: changes in science, change in science]. Almagest: International Journal for the History of Scientific Ideas, 14(2):290–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.136774.

Wagner:2023:CCV

[263] Pierre Wagner. Conceptual change vs disagreement in science. Almagest: International Journal for the History of Scientific Ideas, 14 (2):300–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.136775.

Anonymous:2023:GA

[264] Anonymous. Guidelines for authors. Almagest: International Journal for the History of Scientific Ideas, 14 (2):311–??, ???? 2023. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMA.5.137947.

Anonymous:2024:FMTa

[265] Anonymous. Front matter (table of contents). Almagest: International Journal for the History of Scientific Ideas, 15(1):1–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.nethttps://
                                        www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.142916.

Szczeciniarz:2024:VLP

[266] Jean-Jacques Szczeciniarz. Vuillemin et Lagrange: une philosophie pour les mathématiques. (French) [Vuillemin and Lagrange: a philosophy for mathematics]. Almagest: International Journal for the History of Scientific Ideas, 15(1):6–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.nethttps://
www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.141827.

Malpangotto:2024:RAC

[267] Michela Malpangotto. Une révolution avant Copernic. l’astronomie précopernicienne issue des Theoricæ novæ Georg Peurbach. (French) [A revolution before Copernicus. Pre-Copernican astronomy from the Theoricæ novæ of Georg Peurbach]. Almagest: International Journal for the History of Scientific Ideas, 15 (1):40–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.nethttps://
www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.141828.

Savoie:2024:PCS

[268] Denis Savoie. Le pseudo-cadran solaire de Sainte Irène d’Istanbul. (French) [The pseudo-sundial of Saint Irene of Istanbul]. Almagest: International Journal for the History of Scientific Ideas, 15(1):58–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.nethttps://
www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.141829.

Chahi:2024:GPD

[269] Azeddine Chahi. The geocentric program during presocratic times. Almagest: International Journal for the History of Scientific Ideas, 15 (1):74–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.nethttps://
                                        www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.141830.

Anonymous:2024:BRa

[270] Anonymous.  Book reviews.  Al-

magest: International Journal for the History of Scientific Ideas, 15 (1):94–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.nethttps://
                                        www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.141831.

Anonymous:2024:BMGa

[271] Anonymous. Back matter (guidelines to the authors). Almagest: International Journal for the History of Scientific Ideas, 15(1):109–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.nethttps://
                                        www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.142919.

Anonymous:2024:FMTb

[272] Anonymous. Front matter (“Table of contents”). Almagest: International Journal for the History of Scientific Ideas, 15(2):1–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.nethttps://
                                        www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.145220.

Guicciardini:2024:TER

[273] Niccolò Guicciardini. Time and the Earth’s rotation in Newton’s Principia. Almagest: International Journal for the History of Scientific Ideas, 15(2):4–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.nethttps://
www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.144209.

Rui:2024:APS

[274] Zhang Rui and Guan Yuzhen. The Aristotelian paradigm of “saving the phenomena” in Geminos’s Introduction to the Phenomena. Almagest: International Journal for the History of Scientific Ideas, 15(2):30– ??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.nethttps://
www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.144210.

Megremi:2024:ELI

[275] Athanasia Megremi. Epigrammatic literature: an invaluable addition to the textual tradition of early medieval mathematics or poetry as mathematical text. Almagest: International Journal for the History of Scientific Ideas, 15(2):50–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.nethttps://
www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.144211.

Andriotis:2024:BIT

[276] Antonis N. Andriotis. Out of the blue: An ignored theory of electricity proposed by Benjamin Lesvios. Almagest: International Journal for the History of Scientific Ideas, 15 (2):68–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

//www.brepolsonline.nethttps://
www.brepolsonline.net/doi/epdf/
10.1484/J.ALMAGEST.5.144212.

Anonymous:2024:BRb

[277] Anonymous.  Book reviews.  Al-

magest: International Journal for the History of Scientific Ideas, 15 (2):80–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.nethttps://
                                        www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.144213.

Anonymous:2024:BMGb

[278] Anonymous. Back matter (“Guidelines for authors”). Almagest: International Journal for the History of Scientific Ideas, 15(2):96–??, ???? 2024. CODEN ???? ISSN 1792-2593. URL https:

                                        //www.brepolsonline.nethttps://
                                        www.brepolsonline.net/doi/epdf/
                                        10.1484/J.ALMAGEST.5.145221.